OpenStudy (anonymous):

5X +3Y = 2 and 6X and 4Y=9 I can not get them to equal

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

but for realz i don't understand what you're trying to say.

OpenStudy (anonymous):

just the "and 6X and"

OpenStudy (anonymous):

I have to put numbers in the x and y to get them to equal 2 for the first prob and 9 on the second prob thanks

OpenStudy (anonymous):

I should have put in a + sign for and sorry

OpenStudy (anonymous):

igggghhht so you got: 5x + 3y = 2 and 6x + 4y = 9 well one thing you can do is to get rid of one coefficient, let's say the y one. mutliply the first equation by 4, and the second one by 3, then substract them: 20x + 12y = 8 - 18x + 12y = 27 = 2x = -19 x = -19/2 plug that in ant of the initial equations to find y

OpenStudy (anonymous):

I understand what you did here..thank you very much. My student is saying thank you also. How do you do it for Y? Plug in the X answer for Y and solve?

OpenStudy (anonymous):

exactly, you can plug it in the first or second initial equation and it should give you the same answer

OpenStudy (anonymous):

we get 190 =8 ..does that make sense? 23.75 = y ..does that sound right?

OpenStudy (anonymous):

we get two different answers, 23.75 and 171

OpenStudy (anonymous):

oops she got 23.75 for first equation and 6.3 for the second equation

OpenStudy (anonymous):

ok well if we have x = -19/2: 5x + 3y =2 -95/2 + 3y = 2 y = (2 + 95/2)/3 = 16.5 6x + 4y = 9 -57 + 4y = 9 4y = 66 y = 66/4 = 16.5

OpenStudy (anonymous):

You are really good. Mariah says thanks and what about 2x +5y =16 and 4x +7y=23